# Catalog of Mock Stellar Streams in Milky Way-like Galaxies

### v1

@@ 5 Conclusions
  • We have generated a catalog of GC stellar streams in four MW-like galaxies within a cosmological context of the TNG and FIRE simulations. The catalog combines the Chen & Gnedin (2024a) GC evolution model with the Chen et al. (2025b) algorithm to generate streams in the time-evolving potentials. The catalog contains the full phase-space information for individual stars in each stream, as well as orbital history and physical properties for stream progenitors.
@@
  • We apply mock photometry to our streams and quantify the number of observable stars per stream with Gaia, LSST, and Roman. We find that both LSST and Roman should discover hundreds of new streams in the Milky Way. In particular, Roman will be able to observe stars down to $0.1\,{\mathrm{M}_{\odot}}$ for the majority of streams.
@@ Acknowledgments
agama (Vasiliev, 2019), astropy (Astropy Collaboration et al., 2022), isochrones (Choi et al., 2016), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020)
Appendix A Calibration of the Galactic Potential
Figure 11: Convergence of the halo and disk potentials for TNG galaxy 523889. For each data point, we increment $l_{\rm max}$ or $m_{\rm max}$ and plot the average absolute relative difference across the initial cluster locations compared to the previous point. The solid lines show the convergence using the first TNG snapshot used in our model and the dotted lines shows the last snapshot.
@@ Appendix A Calibration of the Galactic Potential
For both the halo and disk potentials, we sample coefficients $\Phi_{l,m}(r)$ and $\Phi_{m}(R,z)$ on logarithmically spaced grids and spline interpolate between the individual points to get the final analytic potential at each snapshot. The time-dependent potential is then constructed by linearly interpolating between the individual snapshot potentials. For the multipole expansion, no symmetry is assumed and all $m$ terms are kept for each corresponding $l$. To determine the highest order of $l$ in the multipole expansion and $m$ in the Fourier harmonic expansion, we perform a convergence test by evaluating the change in each respective potential at the locations of the GCs 3.5 Gyr ago for one of our halos (523889). For the disk potential, we start at $m_{\rm max}=1$ and compute the relative absolute change in potential between $m_{\rm max}-1$ and $m_{\rm max}$, averaged at the locations of all the GCs. The same is done for the halo potential, except we increment $l_{\max}$ by 2 instead of 1. The result is shown in Figure 11. The shorter lines are for $m_{\rm max}$ and the longer ones are for $l_{\rm max}$. Additionally, the solid lines are for the first snapshot used for the BFEs ($z=0.3$), and the dashed lines are for the last snapshot ($z=0$). Based on the convergence of both the disk and halo potentials, we adopt $l_{\rm max}=12$ and $m_{\rm max}=6$.
@@
Appendix B Comparison of Low and High Time Resolution for Mass-Loss Calculation
@@
The fitting of the BFE potentials allows us to calculate the mass-loss rate along the GC orbits at higher resolution than is available with the GC evolution model in Chen & Gnedin (2024a). In Figure 12, we show our computed effective tidal frequency (defined from the eigenvalues of the tidal tensor), for a selected cluster in the catalog on a similar orbit to Pal 5. Each peak in $\Omega_{\rm tid}$ is from a pericenter passage of the GC on its orbit. By comparison, the Chen & Gnedin (2024a) model only computes $\Omega_{\rm tid}$ at full TNG snapshots, the timestamps of which are shown with vertical dashed lines. For our higher resolution calculation using BFE, agama interpolates the computed solution to give us the values of $\Omega_{\rm tid}$ on $10^{3}$ uniformly space intervals within the last $\sim$3.5 Gyr. Our higher resolution calculation is used to better capture the mass-loss of the GCs along their orbits.
Appendix C Properties of Surviving GC Population
For the creation of our GC stream catalogs, we modified the GC evolution model in Chen & Gnedin (2024a) to output cluster masses at each snapshot and subsequently had to re-run the model on the four cosmological halos. Due to inherent random number generation in the code, our GC populations are slightly different than the published catalogs in Chen & Gnedin (2024a). Here we show the properties of each GC population in comparison with the MW. Figure 13 shows the cumulative probability distribution functions (CDFs) for the mass, metallicity, and radial distributions of surviving GCs in our catalog, with masses greater than $10^{4}\,{\mathrm{M}_{\odot}}$ in the present day. Observed CDFs for the MW GC population are displayed with a thicker line. Our GC populations are matched well particularly to the mass and radial distributions of the MW GCs, both of which are more important for the purposes of generating stellar streams.

### v2

@@ 5 Conclusions
  • We extend the work of Pearson et al. (2024) by generating a catalog of GC stellar streams in four MW-like galaxies within a cosmological context of the TNG and FIRE simulations. The catalog combines the Chen and Gnedin (2024a) GC evolution model with the Chen et al. (2025b) algorithm to generate streams in the time-evolving potentials. The catalog contains the full phase-space information for individual stars in each stream, as well as orbital history and physical properties for stream progenitors.
@@
  • We apply mock photometry to our streams and quantify the number of observable stars per stream with Gaia, LSST, and Roman. We find that both LSST and Roman may discover up to hundreds of new streams in the Milky Way. In particular, Roman will be able to observe stars down to $0.1\,{\mathrm{M}_{\odot}}$ for the majority of streams.
@@ Acknowledgments
agama (Vasiliev, 2019), astropy (Astropy Collaboration et al., 2022), isochrones (Morton, 2015), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020)
Appendix A Properties of Surviving GC Population
For the creation of our GC stream catalogs, we modified the GC evolution model in Chen and Gnedin (2024a) to output cluster masses at each snapshot and subsequently had to re-run the model on the four cosmological halos. Due to inherent random number generation in the code, our GC populations are slightly different than the published catalogs in Chen and Gnedin (2024a). Here we show the properties of each GC population in comparison with the MW [^4] The data used to calculate the MW CDFs is available at https://people.smp.uq.edu.au/HolgerBaumgardt/globular/. Figure 13 shows the cumulative probability distribution functions (CDFs) for the mass, metallicity, and radial distributions of surviving GCs in our catalog, with masses greater than $10^{4}\,{\mathrm{M}_{\odot}}$ in the present day. Observed CDFs for the MW GC population are displayed with a thicker line. Our GC populations are matched well particularly to the mass and radial distributions of the MW GCs, both of which are more important for the purposes of generating stellar streams.
Appendix B Calibration of the Galactic Potential
Figure 11: Convergence of the halo and disk potentials for TNG galaxy 523889. For each data point, we increment $l_{\rm max}$ or $m_{\rm max}$ and plot the average absolute relative difference in the value of the potential across the initial cluster locations compared to the previous point. The solid lines show the convergence using the first TNG snapshot used in our model and the dotted lines shows the last snapshot.
@@ Appendix B Calibration of the Galactic Potential
For both the halo and disk potentials, we sample coefficients $\Phi_{l,m}(r)$ and $\Phi_{m}(R,z)$ on logarithmically spaced grids and spline interpolate between the individual points to get the final analytic potential at each snapshot. The time-dependent potential is then constructed by linearly interpolating between the individual snapshot potentials. For the multipole expansion, no symmetry is assumed and all $m$ terms are kept for each corresponding $l$. To determine the highest order of $l$ in the multipole expansion and $m$ in the Fourier harmonic expansion, we perform a convergence test by evaluating the change in each respective potential at the locations of the GCs 3.5 Gyr ago for one of our halos (523889). For the disk potential, we start at $m_{\rm max}=1$ and compute the relative absolute change in potential between $m_{\rm max}-1$ and $m_{\rm max}$, averaged at the locations of all the GCs. The same is done for the halo potential, except we increment $l_{\max}$ by 2 instead of 1. The result is shown in Figure 11. The shorter lines are for $m_{\rm max}$ and the longer ones are for $l_{\rm max}$. Additionally, the solid lines are for the first snapshot used for the BFEs ($z=0.3$), and the dashed lines are for the last snapshot ($z=0$). Below $\Delta\Phi/\Phi$ of $10^{-3}$, the error in the orbit will be dominated by the resolution of the snapshot interpolation rather than the fits to the potentials themselves. Based on the convergence of both the disk and halo potentials, we adopt $l_{\rm max}=12$ and $m_{\rm max}=6$.
Appendix C Reconstructed Orbits
In Figure 14, we show some reconstructed orbits from our fitted potentials compared to the true particle orbit in the respective cosmological simulation. For each individual halo we show two reconstructed GC orbits for a total of eight. In each panel, the blue dots show the galactocentric radius of the GC (more specifically, the galactocentric radius of the particle in the simulation we assign to represent the GC) at each snapshot used in this work. The top row shows four orbits from the two TNG50 galaxies, and the bottom row shows orbits from the FIRE galaxies. For FIRE, we opt to show the last $\sim 500$ Myr of the orbit, which contains 10 snapshots with high resolution. For TNG50 we show the full integrated orbit used in this work, approximately $3.5$ Gyr for each halo. Our method reliably recovers the pericenter and apocenter values for most orbits.
Figure 14: Reconstructed orbits from our method outlined in Section 2. The blue points represent the true orbit from the cosmological simulation; the black line is our reconstructed orbit calculated from agama using our BFE potentials. The background cosmological simulation for each particular orbit is shown in the top right of each panel.
@@ Appendix C Reconstructed Orbits
Appendix D Comparison of Low and High Time Resolution for Mass-Loss Calculation
@@
The fitting of the BFE potentials allows us to calculate the mass-loss rate along the GC orbits at higher resolution than is available with the GC evolution model in Chen and Gnedin (2024a). In Figure 12, we show our computed effective tidal frequency (defined from the eigenvalues of the tidal tensor), for a selected cluster in the catalog on a similar orbit to Pal 5. Each peak in $\Omega_{\rm tid}$ is from a pericenter passage of the GC on its orbit. By comparison, the Chen and Gnedin (2024a) model only computes $\Omega_{\rm tid}$ at full TNG snapshots, the timestamps of which are shown with vertical dashed lines. For our higher resolution calculation using BFE, agama interpolates the computed solution to give us the values of $\Omega_{\rm tid}$ on $10^{3}$ uniformly space intervals within the last $\sim$3.5 Gyr. Our higher resolution calculation is used to better capture the mass-loss of the GCs along their orbits.
Appendix E Comparison with N-body
Here, we perform an N-body realization of one stream from our catalog to assess the validity of the Chen et al. (2025b) algorithm in a time-evolving, asymmetric halo. For our N-body test we use the multipole gravity solver falcON (Dehnen, 2000, 2002) with our TNG potential from halo 523889 as the external field. We select stream 29 from the 523889 catalog as our test case, which has a pericenter of $\sim$11 kpc and an apocenter of $\sim$21 kpc, similar to the stellar stream Pal 5. We initialize the cluster as a Plummer profile with a scale radius of $4$ pc and initial mass $2.4\times 10^{4}\,{\mathrm{M}_{\odot}}$, which are the same initial conditions provided to the Plummer potential in the catalog realization of the stream. We take the particle masses to be the mean particle mass of stars in the catalog stream, which in this case is $0.33\,{\mathrm{M}_{\odot}}$, making the total number of particles $7.3\times 10^{4}$. We place the N-body cluster at the same initial location in the galaxy and evolve it for $3.5$ Gyr. Based on the optimal softening length relation from Athanassoula et al. (2000), we take $\varepsilon=0.2$ pc and use a timestep of $2^{-13}\penalty 10000\ \mathrm{km\penalty 10000\ s^{-1}\penalty 10000\ kpc^{-1}}\approx 0.1\penalty 10000\ \mathrm{Myr}$.
Figure 15: N-body comparison of one stream from the catalog. In each panel, the top stream is generated from the Chen et al. (2025b) particle-spray algorithm, while the bottom stream is generated using falcON. For clarity of display, the streams are offset by $5^{\circ},60\penalty 10000\ \mathrm{km\penalty 10000\ s^{-1}}$, and $1\penalty 10000\ \mathrm{mas\penalty 10000\ yr^{-1}}$, respectively. The streams are nearly indistinguishable in their distributions along $\phi_{1}$, indicating that the Chen et al. (2025b) algorithm is a valid approximation for full N-body integration even in a time-evolving, non-axisymmetric halo.
Figure 15 compares the positions, proper motions, and radial velocity of the stream generated with falcON compared to the catalog stream generated with particle-spray. The progenitor locations for each stream are approximately located at $\phi_{1}\approx 0^{\circ}$. For plotting purposes, we do not show the progenitor from the N-body stream since we do not resolve the progenitor in the particle-spray version. Additionally, the particle-spray stream has more particles than the final N-body stream, so we randomly down-sample the particle-spray stream to show the same number of particles in both streams. Finally, we apply offsets of $(5^{\circ},60\penalty 10000\ \mathrm{km\penalty 10000\ s^{-1}},1\penalty 10000\ \mathrm{mas\penalty 10000\ yr^{-1}},1\penalty 10000\ \mathrm{mas\penalty 10000\ yr^{-1}})$ to the particle-spray stream on the y-axis so that the two streams can be visually compared. Comparison of the tracks along $\phi_{1}$ shows that the streams are nearly indistinguishable in morphology, proper motions, and radial velocity, with good agreement for both the overall stream length and width. Using the same classification statistics from Figure 5, the percent differences between the particle-spray stream and N-body stream $(\sigma_{\phi_{1}},\sigma_{\phi_{2}},\sigma_{\mu_{\phi_{1}}},\sigma_{\mu_{\phi_{2}}})$ are $(-13\%,-26\%,-12\%,-12\%)$. The difference in the radial velocity spread, $\sigma_{rv}$ is $6\%$.
Across these metrics, the particle-spray model tends to have a slightly smaller dispersion than the N-body model. In addition to inherent errors caused by differences between the two methods (N-body versus particle-spray), our prescription for GC mass loss may also contribute to the difference in some quantities. For example, part of the difference in $\sigma_{\phi_{1}}$ could be attributed to the number of stars unbound per unit time according to Equation 2 versus the value from direct N-body. We also find that different choices for the softening length have an effect on how well the resulting stream matches the particle-spray; for example, a trial run with the same setup but a softening length of $\varepsilon=1\penalty 10000\ \mathrm{pc}$ produces better agreement with the particle-spray model, with percent differences in $(\sigma_{\phi_{1},}\sigma_{\phi_{2}},\sigma_{\mu_{\phi_{1}}},\sigma_{\mu_{\phi_{2}}},\sigma_{rv})$ being (-6%, -17%, -3%, -1%, +9%). However, given the size of our initial cluster and the number of particles used, this softening length is somewhat large to accurately capture the internal cluster dynamics.
Aside from the width, the deviation of all other metrics are still around $10\%$, which is the average value quoted in Chen et al. (2025b) for the potential models tested in their work. Based on this as well as the visual similarity of the stream tracks, we conclude that the Chen et al. (2025b) algorithm is still a valid approximation for modeling cold streams in a time-varying, non-axisymmetric halo.